OpenStudy (anonymous):

The sum of the first 2 terms of a geometric sequence is 8 and the 3rd term of sequence is 18. Find the 1st term of the sequence.

OpenStudy (anonymous):

@ganeshie8 @RadEn @hartnn

OpenStudy (zzr0ck3r):

a+b=8 c=18 b/a=c/b use b=8-a (8-a)/a=18/(8-a) solve this and get a = 2 or 32 its easy to see 32 wont work so a = 2 and thus b = 6

OpenStudy (zzr0ck3r):

1st term = 2

OpenStudy (anonymous):

why 32 is not the answer...

OpenStudy (zzr0ck3r):

because the third term is 18, so that would mean you have negative terms, but the 3rd term is positive, and this is geometric

OpenStudy (anonymous):

geometric sequence can have negative ratio..

OpenStudy (zzr0ck3r):

well ok maybe it can be both? that would make the second term -24 and -24/32 = -3/4 18/-24 = -3/4

OpenStudy (zzr0ck3r):

ok so both will work....

OpenStudy (anonymous):

then 2 and 32 are the answers?

OpenStudy (anonymous):

thx

OpenStudy (zzr0ck3r):

yeah, sort of odd that it asked for one answer